# Writing – review & editing, Writing – original draft, Visualization, Data curation ,Investigation, Formal analysis, Validation, Software, Methodology, Conceptualization

### v1

@@ 3.3 Evaluating the Quality of Adversarial Attacks
The quality of an adversarial attack can be derived from the extent to which the crafted samples maintain coherence and consistency, making them indistinguishable from benign samples. To assess this, we propose two evaluation criteria for measuring the distinguishability of adversarial samples: first, by evaluating how anomalous adversarial samples are compared to benign samples, and second, by assessing the extent to which adversarial samples influence the internal behavior of the target model when queried.
Assessing the anomaly of the adversarial sample. To evaluate the distinguishability of adversarial samples, we propose using an anomaly detection rate metric that measures the percentage of samples identified as anomalies for each attack. This rate can be indicative of how different the adversarial samples are from the benign distribution. We calculated the anomaly detection rate using two established methods from the anomaly detection domain: i) the anomaly scores provided by the isolation forest method (IF) [30]; and ii) the reconstruction error from an autoencoder (AE) [44]. Both the anomaly score and the reconstruction error indicate the likelihood of an adversarial sample remaining coherent and consistent, and thus being indistinguishable from the original benign samples. When these values are low, the sample is considered less anomalous, and therefore less distinguishable as an adversarial sample, and vice versa.
Our proposed methodology evaluates the anomaly rate separately for each target class since an adversarial sample can be anomalous for the target class while still maintaining the overall distribution of the dataset — especially when the changes made to the sample are subtle, such that the sample maintains the same distribution as the original sample but is detected as anomalous when compared to the intended target class.
For each class, an anomaly detection model is trained on benign samples (i.e., the data used to train the target model). Then, based on the target model’s prediction for the adversarial sample, we query the relevant anomaly detection model with the adversarial sample. For example, if the adversarial sample was classified as ‘1’ by the target model, it would be evaluated using the anomaly detection model trained on benign samples from class 1.
@@
In the first method, a sample is considered anomalous if the IF model identifies it as an outlier. For the second method, a sample is considered anomalous if its reconstruction error, calculated by the AE, exceeds a predefined threshold; this threshold is defined in Eq. 1:
$Threshold(x,AE)=mean(test_{r.\,error})+0.5*std(test_{r.\,error})$
where $x$ is a data sample, $AE$ is the autoencoder, and $test_{r.\,error}$ is the reconstructed error on the test set when training the AE.
@@
The AE model consists of an embedding layer with 64 neurons, and it was trained using the Adam optimizer with a learning rate of $31e-$ 3, a weight decay of $81e-$ 8, for 10 epochs, and the mean square error (MSE) as the loss. The IF method, in turn, was implemented with 100 estimators, a maximum features set at one, and without bootstrapping. For comparison, the contamination parameter in the IF model was set to match the false positive rate (FPR) obtained by the AE according to the threshold defined in Eq. 1, corresponding to the proportion of benign samples with reconstruction error above this threshold.
Assessing the impact of adversarial samples on the target model.
An additional aspect of adversarial sample quality is whether they introduce inconsistency in the target model’s decision-making process. Adversarial samples can affect the target model’s decision by altering feature importance.
To quantify this impact, we developed new criteria using SHAP values [32], a widely used XAI technique that provides a unified measure of feature importance in ML models. SHAP values estimate the contribution of each feature to the prediction by comparing the actual predictions with averages across all possible feature combinations. we calculated the difference in SHAP values between benign and adversarial samples, focusing on the top-four most important features selected based on their average SHAP values computed based on the benign samples in the training dataset. The SHAP values were extracted using the SHAP TreeExplainer.
@@
We defined two concrete metrics to evaluate this impact: i) the importance-based anomaly detection rate, which measures the percentage of adversarial samples where at least one feature’s SHAP value exceeds the maximum or falls below the minimum SHAP value for that feature in the benign dataset; and ii) the average number of features with anomalous SHAP values per sample. These metrics reflect the extent to which adversarial perturbations influenced the target model. A lower number of features with anomalous SHAP values and a reduced importance-based anomaly detection rate indicate higher quality adversarial samples, as the model’s internal behavior remains more consistent despite changes in predictions.
In addition to serving as a metric assessing the quality of adversarial samples, the anomaly detection rate metrics, whether computed on adversarial samples or SHAP values, can serve as the basis of a defense mechanism, as they are strong indicator of attacked samples.
@@ 4.1 Research Questions
  • RQ#2 How the perturbed samples of different attacks’ strategies differ in terms of the attacker’s risk and effort?
@@ 4.5 Attack Configuration
In our comparison of different types of attacks, we used two query-based attacks and five transferability-based gradient attacks that varied in terms of their feature selection techniques. The boundary attack was used with an epsilon of one, a delta of one, maximal iterations of 3000, the number of trials set at 20, and an adaptation step of one. The HopSkipJump attack was used with the L2 norm, maximal iterations of 50, the max-eval parameter set at 10000, the init-eval set at 500, and an init size of 100. The parameters of the boundary and HopSkipJump attacks were identical for all datasets, except for the maximal iteration number; for the VideoTQ dataset this was set at 1000 in the HopSkipJump attack.
@@
The transferability-based attacks require the attacker to train a surrogate model (illustrated in Fig. 1 (d, e)). In our evaluation, all surrogate models were NNs, and we used an embedding layer to try to maintain the consistency of the crafted samples, following the approach employed in other studies [35, 19]. All surrogate models were comprised of two components: a sub-model for embedding and a sub-model for classification, both of which were fine-tuned for optimized performance. The embedding sub-model had an input size matching the number of features in the input sample, followed by a dense layer with 256 neurons, and ReLU and PReLU activations for the Hate and VideoTQ datasets respectively. For the ICU dataset, the embedding sub-model had three dense layers with 256, 128, and 64 neurons and ReLU activation for all layers. The output size of all embedding sub-models was set at 16, i.e., the embedding size. The classification sub-model had a dense layer with 16 neurons (matching the embedding size) with ReLU activation for the Hate and ICU datasets, and PReLU activation for the VideoTQ datasets. This was followed by a dropout layer with a dropout rate of 0.1. All surrogate models were trained using the binary cross-entropy loss function [31] and the Adam optimizer [26] with a learning rate of 0.2 for the Hate and VideoTQ datasets, and 0.1 for the ICU dataset.
After training the surrogate models, the attacker selects $k$ features and identifies the $n$ features most correlated with them, which are then perturbed during each attack iteration. Each transferability-based attack uses a different selection technique to determine which features to perturb (see Appendix A). In attacks employing importance-based selection techniques, features are selected based on their importance as determined by XGBoost, GradientBoost, LightGBM, and Random Forest models, trained separately for these attacks. For all attacks, $n$ was set at one, and the correlated features were selected by using Pearson’s correlation coefficient [48], where $k$ was set at two;
@@ 4.6 Regression Models
Of the datasets used in our evaluation, the ICU dataset is the only one containing dependent features where their dependencies are not well-defined. Specifically, these dependent features are: ‘apache_3j_bodysystem,’ ‘apache_3j_diagnosis,’ ‘d1_mbp_invasive_max,’ and ‘d1_mbp_invasive_min.’
@@ 4.7 Threat Model
We assume that the attacker can query the target model and has a query budget of one for transferability-based attacks and an unlimited query budget for query-based attacks; in each query to the target model, the attacker obtains the confidence score of the prediction. In addition, we assume that the attacker has no prior knowledge about the specific model architecture or the ML algorithm used. The attacker also has no access to the internal parameters of the model (such as weights and biases) and cannot calculate any gradients related to the model. Furthermore, the attacker does not have access to the target model’s training data; However, we assume that the attacker has a surrogate dataset derived from the same distribution, meaning that it contains the same features, in the same order, as the training data.
@@ 4.8 Evaluation Setup
For a robust evaluation, we filtered the test set by selecting only samples correctly classified by both the target and surrogate models, excluding those already misclassified, retaining $85\%$, $69\%$, and $91\%$ of samples from the Hate, ICU, and VideoTQ datasets respectively. In addition, to properly evaluate each attack, we randomly selected a balanced number of samples from each class, which were then used as the attack set. This resulted in attack sets that contained $182$, $1000$, and $1000$ samples for the Hate, ICU, and VideoTQ datasets respectively. In the evaluation of each adversarial attack’s samples, only those samples that successfully misled both the surrogate and target models (or only the target model in the case of a query-based attack) were used.
@@ 4.9 Evaluation Metrics
The main objective of our evaluation is to analyze the characteristics of and differences between black-box query- and transferability-based adversarial attacks with respect to three key factors: the attacker’s risk, the attacker’s effort, and the attacks’ quality (illustrated in Fig. 1(f)). To perform a thorough analysis, we used the evaluation metrics commonly used in the literature and the metrics proposed in this work.
@@
The attacks’ quality is reflected in the crafted adversarial sample, which were assessed using three metrics (illustrated in Fig. 1(f-3)): i) the anomaly detection rate of adversarial samples compared to benign samples, using IF and AE models, where a higher rate indicates less consistency and coherence; ii) the percentage of samples with at least one feature with an anomalous SHAP value (importance-based anomaly detection rate); and iii) the average number of features with anomalous SHAP values per sample.
@@ 5.2 Attacker’s Risk
Table 2 provides details on the attack’s success rate (SR) on the target model, the attack’s success rate on the surrogate model (Surrogate SR), the transfer success rate from the surrogate model to the target model (Transfer SR), and the overall success rate (Overall SR), which represents the percentage of adversarial samples that successfully mislead both the surrogate model and the target model.
Table 2: Attacker’s risk: performance of query- and transferability-based attacks on different target models, including the attack’s success rate (SR) on the target model, the attack’s success rate on the surrogate model (Surrogate SR), the transfer success rate from the surrogate model to the target model (Transfer SR), and the overall success rate (Overall SR), which represents the percentage of adversarial samples that successfully mislead both the surrogate model and the target model (imp. refers to an importance-based feature selection technique).
@@ 5.3 Attacker’s Effort
With respect to the amount of data required, the attacker’s objective is to minimize the amount of data. In query-based attacks, the only data required by the attacker is the original sample that the adversarial sample will be based on, however in transferability-based attacks, the attacker must also possess a surrogate dataset to train the surrogate and feature importance models.
Based on this, we can conclude that transferability-based attacks require more data samples than query-based attacks.
@@ 5.4 Attack’s Quality
Figure 3: Attacks’ quality evaluated based on the anomaly detection rate by AE and IF models across datasets, computed separately for each class. The FPR on all datasets for both classes is less than 0.04.
Coherence of the adversarial sample.  Fig. 3 presents the anomaly detection rates derived from the values obtained using the IF model and the reconstruction errors from the AE for adversarial samples for each evaluated attack, target class, and dataset. Each anomaly detection rate represents the percentage of adversarial samples identified as anomalies. For the AE, a sample is considered anomalous if its reconstruction error exceeds a predefined threshold (see  Eq. 1. The false positive rate (FPR) for benign samples, based on the specified threshold, is 0.020 and 0.011 for class 0 and class 1 on the Hate dataset, 0.038 and 0.038 on the ICU dataset, and 0.014 and 0.002 on the VideoTQ dataset. The same FPR was applied to the IF model for comparison.
@@
Impact on the target model’s decision-making process.
Table 4: Attacks’ quality: results based on the importance-based anomaly detection rate (percentage of samples with at least one feature with an anomalous SHAP value) and the average number of anomalous features computed based on their SHAP values. Bold values highlight samples with the fewest changes in SHAP values, indicating high-quality samples in terms of having minimal impact on the target model’s decision process.
@@
Importance-based Anomaly Detection Rate (%) / Avg. Number of Anomalous Features per Sample
Hate GB 100.0 / 17.9 100.0 / 16.2 100.0 / 23.4 100.0 / 22.4 81.2 / 8.6 32.7 / 0.7 44.9 / 1.2
LGB 69.8 / 1.8 65.9 / 2.0 92.9 / 4.1 98.4 / 4.4 95.1 / 3.0 71.4 / 1.4 53.6 / 0.9
XGB 45.1 / 0.9 35.7 / 0.5 100.0 / 6.6 100.0 / 6.2 80.3 / 3.2 13.8 / 0.2 36.4 / 0.5
RF 85.2 / 2.1 87.4 / 1.8 100.0 / 8.6 100.0 / 9.2 78.4 / 6.0 64.3 / 1.2 68.9 / 1.4
ICU GB 83.9 / 1.9 74.9 / 1.6 80.5 / 1.3 71.4 / 1.3 14.0 / 0.2 17.3 / 0.2 20.7 / 0.3
LGB 17.1 / 0.2 25.9 / 0.3 77.7 / 1.6 67.8 / 1.4 22.4 / 0.3 31.4 / 0.4 24.8 / 0.3
XGB 23.3 / 0.3 23.0 / 0.3 48.1 / 0.7 46.2 / 0.6 16.3 / 0.2 15.9 / 0.2 14.9 / 0.2
RF 10.5 / 0.1 9.8 / 0.1 9.0 / 0.1 10.2 / 0.1 11.2 / 0.1 10.3 / 0.1 12.0 / 0.1
VideoTQ GB 24.1 / 0.3 11.8 / 0.6 1.2 / 0.0 2.6 / 0.1 66.7 / 2.7 100.0 / 4.0 71.5 / 3.2
LGB 29.0 / 0.3 28.9 / 0.3 1.8 / 0.0 2.0 / 0.0 4.3 / 0.0 10.3 / 0.1 7.4 / 0.1
XGB 70.9 / 1.6 71.7 / 1.1 12.6 / 0.1 13.3 / 0.2 49.3 / 0.6 80.6 / 1.3 62.7 / 0.9
RF 73.7 / 1.3 95.2 / 1.7 1.8 / 0.0 2.0 / 0.0 41.2 / 0.4 92.2 / 1.3 93.4 / 1.5
@@
Figure 4: Attacks’ quality evaluated based on the impact on the target model’s decision-making process; SHAP value distribution for benign (blue) and adversarial samples (on the Hate dataset), across the top-four most important features selected based on their average SHAP values computed on the benign training dataset.
@@
Figure 5: Attacks’ quality evaluated based on the impact on the target model’s decision-making process; SHAP value distribution for benign (blue) and adversarial samples (on the ICU dataset), across the top-four most important features selected based on their average SHAP values computed on the benign training dataset.
@@
Figure 6: Attacks’ quality evaluated based on the impact on the target model’s decision-making process; SHAP value distribution for benign (blue) and adversarial samples (on the VideoTQ dataset), across the top-four most important features selected based on their average SHAP values computed on the benign training dataset.

### v2

@@ 3.3 Evaluating the Quality of Adversarial Attacks
The quality of an adversarial attack is measured by its ability to evade detection, which depends on how well the crafted samples maintain coherence and consistency with benign data. To assess this, we propose two evaluation criteria:
  1. 1.
    Feature Space Coherency – Quantifying the distinguishability of adversarial samples from benign data distributions.
  2. 2.
    Model Interpretation Stability – Analyzing the extent to which adversarial samples influence the internal decision-making process of the target model when queried.
Both criteria leverage a key contribution of our work: the introduction of Class-Specific Anomaly Detection (CSAD). Unlike traditional one-class anomaly detection methods that assume all benign samples belong to a single class or distribution, CSAD evaluates anomalies relative to the specific class to which the adversarial sample is assigned by the target model. This class-aware evaluation enables CSAD to account for class-dependent variations in benign data, resulting in more effective anomaly detection.
The CSAD Approach. In the CSAD approach, we evaluate anomalies separately for each target class. For instance, given a dataset with three classes, we train three distinct anomaly detection models—one for each class. Each model is trained exclusively on benign samples from its respective class, ensuring that the unique statistical properties of each class-specific distribution are captured. When an adversarial sample is generated, it is evaluated using the anomaly detection model corresponding to its predicted class. For example, if an adversarial sample is classified as class 1 by the target model, it is assessed using the anomaly detection model trained on benign samples from class 1.
This approach addresses a key limitation of traditional anomaly detection techniques, where adversarial samples appear anomalous within the general data distribution, but not within their assigned class. Adversarial attacks introduce subtle perturbations that might preserve the statistical characteristics of the original sample while misleading the model. By design, adversarial samples can still resemble their true class and remain indistinguishable from other benign samples in that class. However, when analyzed in the context of the target class (the class to which they are misclassified), these perturbations may exhibit distinct anomalies. By leveraging class-specific distributions, CSAD enables more precise differentiation between adversarial and benign samples, capturing anomalies that would otherwise go undetected.
Feature Space Coherency. To evaluate the distinguishability of adversarial samples, we propose measuring the anomaly detection rate-the percentage of samples identified as anomalies for each attack. The anomaly detection rate metric serves as an indicator of how different adversarial samples are from the benign data distribution. We compute the anomaly detection rate using two well-established methods from the anomaly detection domain: i) Isolation Forest (IF) – A tree-based algorithm that identifies anomalies based on how easily a sample can be isolated in a feature space, and ii) Autoencoder (AE) – A neural network-based model that reconstructs benign samples and detects anomalies based on the reconstruction error.
@@
In the IF algorithm, a sample is considered anomalous if it is classified as an outlier. In the AE model, an anomaly is identified when its reconstruction error exceeds a threshold, defined in Eq. 1:
$\begin{split}Threshold(x_{c},\>AE_{c})=mean(test_{c,\>r.\,error})\,+\\ 0.5*std(test_{c,\>r.\,error})\end{split}$ (1)
where $x_{c}$ is a sample classified as class $c$, $AE_{c}$ is the autoencoder trained on benign samples of class $c$, and $test_{c,r.,error}$ represents the reconstruction error on a validation set of class $c$ during the AE’s training.
@@
The AE model consists of an embedding layer with 64 neurons, and it was trained using the Adam optimizer with a learning rate of $31e-$ 3, a weight decay of $81e-$ 8, for 10 epochs, and the mean square error (MSE) as the loss. The IF method was implemented with 100 estimators and $max\_features$ set to one, without bootstrapping. For comparison, the contamination parameter in the IF model was set to match the false positive rate (FPR) obtained by the AE according to the threshold defined in Eq. 1, corresponding to the proportion of benign samples with a reconstruction error above this threshold.
Other anomaly detection methods can be considered depending on the dataset characteristics and attack patterns.
Model Interpretation Stability. In addition to measuring distinguishability, we assess how adversarial samples influence the target model’s decision-making process. Adversarial perturbations may alter the model’s perception of feature importance, leading to inconsistencies in the model’s reasoning. To quantify this effect, we employ SHAP [33], a widely used explainability technique that provides a unified measure of feature importance in ML models. SHAP estimates each feature’s contribution to a prediction by comparing actual predictions with averages across all possible feature combinations.
@@
We introduce two metrics based on SHAP: (1) Importance-Based Anomaly Detection Rate – The percentage of adversarial samples where at least one feature’s SHAP value falls outside the range observed in benign samples, and (2) Average Anomalous SHAP Features per Sample – The average number of features in an adversarial sample with SHAP values outside the benign distribution. Lower values for these metrics indicate higher-quality adversarial samples, as they remain consistent with the model’s internal reasoning despite altering predictions.
Importantly, CSAD is also being extended to SHAP-based anomaly detection; since we assess traditional anomaly detection methods per class, we compute SHAP-based anomalies within the context of each class. This ensures that SHAP value deviations are analyzed relative to the specific class distribution, rather than across the entire dataset. The SHAP values were extracted using SHAP TreeExplainer.
Leveraging Anomaly Detection for Defense. The Importance-Based Anomaly Detection Rate and Average Anomalous SHAP Features per Sample methodologies not only assess adversarial quality but also serve as defense mechanisms. Since adversarial samples often exhibit statistical or interpretability inconsistencies, the anomaly detection rate—whether computed directly or through SHAP values—serves as a robust indicator for detecting adversarial attacks. Our class-specific approach further enhances the ability to distinguish between benign and adversarial samples, offering a stronger defense against adversarial manipulation.
@@ 4.1 Research Questions
  • RQ#2 How the perturbed samples of different attack strategies differ in terms of the attacker’s risk and effort?
@@ 4.5 Attack Configuration
In our comparison of different types of attacks, we used two query-based attacks and five transferability-based gradient attacks that varied in terms of their feature selection techniques. The boundary attack was performed with an epsilon of one, a delta of one, maximal iterations of 3000, the number of trials set at 20, and an adaptation step of one. The HopSkipJump attack was performed with the L2 norm, maximal iterations of 50, the max-eval parameter set at 10,000, the init-eval set at 500, and an init size of 100. The parameters of the boundary and HopSkipJump attacks were identical for all datasets, except for the maximal iteration number; for the VideoTQ dataset this was set at 1000 in the HopSkipJump attack.
@@
The transferability-based attacks require the attacker to train a surrogate model (illustrated in Fig. 1 (d, e)). In our evaluation, all surrogate models were neural networks, and we used an embedding layer to try to maintain the consistency of the crafted samples, following the approach employed in other studies [36, 21]. All surrogate models were comprised of two components: a sub-model for embedding and a sub-model for classification, both of which were fine-tuned for optimized performance. The embedding sub-model had an input size matching the number of features in the input sample, followed by a dense layer with 256 neurons and ReLU and PReLU activations for the Hate and VideoTQ datasets respectively. For the ICU dataset, the embedding sub-model had three dense layers with 256, 128, and 64 neurons and ReLU activation for all layers. The output size of all embedding sub-models was set at 16, i.e., the embedding size. The classification sub-model had a dense layer with 16 neurons (matching the embedding size) with ReLU activation for the Hate and ICU datasets and PReLU activation for the VideoTQ datasets. This was followed by a dropout layer with a dropout rate of 0.1. All surrogate models were trained using the binary cross-entropy loss function [32] and the Adam optimizer [28] with a learning rate of 0.2 for the Hate and VideoTQ datasets and 0.1 for the ICU dataset.
After training, the attacker performs the adversarial attack against the surrogate model by perturbing a group of features in each iteration. Each transferability-based attack employs a different feature selection technique to determine which features to perturb (see Appendix A).
For importance-based selection technique, the $k$ most important features are selected, along with the $n$ features most correlated with them—resulting in a total of $k\times n$ additional perturbed features per iteration. In each subsequent iteration, the next $k$ most important features and their $n$ most correlated features are added to the set of perturbed features. This process continues until an adversarial sample is found or no additional features remain for selection. Feature importance is determined using XGBoost, gradient boosting, LightGBM, and random forest models, each trained separately for this purpose. For random-based selection attacks, $k$ features are randomly chosen in each iteration, along with their $n$ most correlated features.
For all attacks, $k$ was set at two and $n$ was set at one (i.e., one correlated feature per selected feature). The correlated features were identified using the Pearson correlation coefficient [49].
@@ 4.6 Regression Models
Of the datasets used in our evaluation, the ICU dataset is the only one containing dependent features whose dependencies are not well-defined. Specifically, these dependent features are: ‘apache_3j_bodysystem,’ ‘apache_3j_diagnosis,’ ‘d1_mbp_invasive_max,’ and ‘d1_mbp_invasive_min.’
@@ 4.7 Threat Model
We assume that the attacker can query the target model and has an unlimited query budget for query-based attacks and a query budget of one for transferability-based attacks; in each query to the target model, the attacker obtains the confidence score for the prediction. In addition, we assume that the attacker has no prior knowledge about the specific model architecture or ML algorithm used. The attacker also has no access to the internal parameters of the model (such as weights and biases) and cannot calculate any gradients related to the model. Furthermore, the attacker does not have access to the target model’s training data; however, we assume that the attacker has a surrogate dataset derived from the same distribution, meaning that it contains the same features, in the same order, as the training data.
@@ 4.8 Evaluation Setup
To ensure a robust evaluation, we created the attack set by filtering the test set to include only samples correctly classified by both the target and surrogate models, excluding those already misclassified. This filtering retained $85\%$, $69\%$, and $91\%$ of samples from the Hate, ICU, and VideoTQ datasets, respectively. Next, we randomly selected a balanced number of samples from each class. As a result, the attack sets contained 182, 1,000, and 1,000 samples for the Hate, ICU, and VideoTQ datasets, respectively. Our analysis of adversarial samples (using $L_{p}$ norms or quality assessment), considered only samples that successfully fooled the target model (or both the surrogate and target models in the case of transferability-based attacks).
@@ 4.9 Evaluation Metrics
The main objective of our evaluation is to analyze the characteristics of and differences between black-box query- and transferability-based adversarial attacks with respect to three key factors: the attacker’s risk, the attacker’s effort, and the attack quality (illustrated in Fig. 1(f)). To perform a thorough analysis, we used the evaluation metrics commonly used in the literature and the metrics proposed in this work.
@@
The attack quality is reflected in the crafted adversarial sample, which were assessed using three metrics (illustrated in Fig. 1(f-3)): i) the anomaly detection rate of adversarial samples compared to benign samples, using IF and AE models, where a higher rate indicates less consistency and coherence; ii) the percentage of samples with at least one feature with an anomalous SHAP value (Importance-Based Anomaly Detection Rate); and iii) the average anomalous SHAP features per sample.
@@ 5.2 Attacker’s Risk
Table 2 presents details on the following attack success rates: Target Success Rate (SR): The percentage of adversarial samples that successfully mislead the target model; Surrogate Success Rate (Surrogate SR): The percentage of adversarial samples that successfully mislead the surrogate model; Transfer Success Rate (Transfer SR): The percentage of adversarial samples that successfully transfer from the surrogate model to the target model, calculated as the ratio of the Surrogate SR. Overall Success Rate (Overall SR): The percentage of adversarial samples that successfully mislead both the surrogate and target models, computed as the ratio of the entire attack set.
Table 2: Attacker’s risk: performance of query- and transferability-based attacks on different target models, evaluated using the following metrics: Target Success Rate (SR): percentage of adversarial samples that successfully mislead the target model; Surrogate Success Rate (Surrogate SR): percentage of adversarial samples that successfully mislead the surrogate model; Transfer Success Rate (Transfer SR): percentage of successful adversarial samples that transfer from the surrogate model to the target model (computed as the ratio of the Surrogate SR); Overall Success Rate (Overall SR): percentage of adversarial samples that mislead both the surrogate and target models (computed based on the entire attack set). Note that importance-based feature selection techniques are referred to as Imp. in the table.
@@ 5.3 Attacker’s Effort
With respect to the amount of data required, the attacker’s objective is to minimize the amount of data. In query-based attacks, the only data required by the attacker is the original sample that the adversarial sample will be based on, however in transferability-based attacks, the attacker must also possess a surrogate dataset to train the surrogate and feature importance models. Based on this, we can conclude that transferability-based attacks require more data samples than query-based attacks.
@@ 5.4 Attack Quality
Figure 3: Attack quality evaluated based on the anomaly detection rate by AE and IF models across datasets, computed separately for each class. The FPR on all datasets for both classes is less than 0.04.
Feature Space Coherency.  Fig. 3 presents the anomaly detection rates derived from the values obtained using the IF model and the reconstruction errors from the AE for adversarial samples for each evaluated attack, target class, and dataset. Each anomaly detection rate represents the percentage of adversarial samples identified as anomalies. For the AE, a sample is considered anomalous if its reconstruction error exceeds a predefined threshold (see  Eq. 1. The false positive rate (FPR) for benign samples, based on the specified threshold, is 0.020 and 0.011 for class 0 and class 1 on the Hate dataset, 0.038 and 0.038 on the ICU dataset, and 0.014 and 0.002 on the VideoTQ dataset. The same FPR was applied to the IF model for comparison.
@@
Model Interpretation Stability.
Table 4: Attack quality: results based on the Importance-Based Anomaly Detection Rate (percentage of samples with at least one feature with an anomalous SHAP value) and the Average Anomalous SHAP Features per Sample computed according to the features’ SHAP values. Bold values are samples with the fewest changes in SHAP values, indicating high-quality samples in terms of having minimal impact on the target model’s decision process.
@@
Importance-Based Anomaly Detection Rate (%) / Average Anomalous SHAP Features per Sample
Hate GB 100.0 / 7.7 96.7 / 5.6 100.0 / 8.4 100.0 / 8.9 81.2 / 3.3 27.3 / 0.4 25.6 / 0.5
LGB 58.2 / 1.1 78.0 / 1.7 64.3 / 1.6 79.7 / 2.0 76.8 / 1.7 69.9 / 1.4 40.5 / 0.7
XGB 26.4 / 0.4 67.6 / 1.1 86.4 / 1.8 92.1 / 1.9 80.3 / 1.6 22.0 / 0.3 24.8 / 0.3
RF 63.2 / 1.0 30.8 / 0.5 100.0 / 3.8 100.0 / 4.2 75.7 / 3.4 67.3 / 0.9 50.0 / 0.6
ICU GB 29.0 / 0.4 19.4 / 0.3 7.3 / 0.1 10.8 / 0.1 1.8 / 0.0 3.5 / 0.0 3.5 / 0.0
LGB 4.1 / 0.0 3.9 / 0.0 22.3 / 0.2 17.5 / 0.2 2.4 / 0.0 6.9 / 0.1 5.6 / 0.1
XGB 3.5 / 0.0 6.0 / 0.1 6.5 / 0.1 6.4 / 0.1 4.2 / 0.0 5.1 / 0.1 3.3 / 0.0
RF 0.1 / 0.0 0.2 / 0.0 4.0 / 0.0 0.5 / 0.0 0.0 / 0.0 0.0 / 0.0 0.7 / 0.0
VideoTQ GB 5.3 / 0.1 9.4 / 0.2 0.4 / 0.0 1.0 / 0.0 51.4 / 0.8 90.3 / 1.5 64.3 / 1.1
LGB 7.3 / 0.1 1.5 / 0.0 3.4 / 0.0 3.4 / 0.0 0.8 / 0.0 0.6 / 0.0 1.0 / 0.0
XGB 38.9 / 0.4 25.2 / 0.3 0.4 / 0.0 1.0 / 0.0 21.6 / 0.2 31.5 / 0.3 17.8 / 0.2
RF 66.6 / 0.7 87.6 / 0.9 0.0 / 0.0 0.0 / 0.0 0.0 / 0.0 68.8 / 0.7 79.0 / 0.8
@@
Figure 4: Attack quality evaluated based on the impact on the target model’s decision-making process; SHAP value distribution for benign (blue) and adversarial samples (on the Hate dataset), across the top-four most important features selected based on their average SHAP values computed on the benign training dataset.
@@
Figure 5: Attack quality evaluated based on the impact on the target model’s decision-making process; SHAP value distribution for benign (blue) and adversarial samples (on the ICU dataset), across the top-four most important features selected based on their average SHAP values computed on the benign training dataset.
@@
Figure 6: Attack quality evaluated based on the impact on the target model’s decision-making process; SHAP value distribution for benign (blue) and adversarial samples (on the VideoTQ dataset), across the top-four most important features selected based on their average SHAP values computed on the benign training dataset.